Night report 2016-01-26

Observing: Nancy Elias-Rosa, Giacomo Terreran, Aleksandar Cikota

Support Team: Stefano Benetti, Lina Tomasella

TAT: Joe Anderson, Morgan Fraser

------------------------------------------------

Weather conditions

Start of night: clear skies, but high humidity (80%)

Middle of night: cloudy, dome closed for ~2.5h

End of night: clear skies

------------------------------------------------

Calibrations

Evening: PESSTO_SOFI_FlatJHK, PESSTO_SOFI_specflat, PESSTO_SOFI_SpecArc

Morning: -

------------------------------------------------

Science

-00:20 UT: Focusing, Image analysis

-00:40 UT: Follow-up: PS15dpn

-01:32 UT: Follow-up: PS15cww

-02:50 UT: Follow-up: OGLE15xl (J and H are good, Ks probably useless because of clouds)

-04:06 UT: Follow-up: ASASSN-14ko

-04:10 UT: ASASSN-14ko ABORTED, dome closed because of clouds

-06:40 UT: dome reopened, clear skies

-06:50 UT: Follow-up: ASASSN-14ko

-07:23 UT: Follow-up: OGLE15xl (Ks band only)

-07:25 UT: OGLE15xl ABORTED because of low altitude/high airmass

-07:26 UT: Standard, Telluric: HIP64944

-07:35 UT: Follow-up: AT2016X

-08:45 UT: Standard, Spectroscopic: EG 274

-08:56 UT: Finished for tonight

------------------------------------------------